OpenStudy (anonymous):

what is the binomial probability of getting 11 girls and 1 boy?

OpenStudy (anonymous):

is this a joke?

Parth (parthkohli):

Nope. Here, you are trying to find the probability of getting 11 boys in 12 "tries". Now, let's apply the binomial probability formula here. Here we go:\[\binom{n}{r}\cdot p^{r} \cdot (1-p)^{n-r}\]

Parth (parthkohli):

So here, \(n = 12, \ r = 11\). We'll assume that the probability of getting a boy or a girl is 0.5. So we are left with this:\[\binom{12}{11}\cdot (0.5)^{11} \cdot (0.5)^{1}\]Simplify and we're done.